OpenStudy (anonymous):

A rectangular prism has a width of 92 ft and a volume of 240 ft3. Find the volume of a similar prism with a width of 46 ft @iGreen

Directrix (directrix):

If two solids are similar, the cube of the scale factor of the two solids is equal to the ratio of the volumes.

Directrix (directrix):

(92/46)^3 = 240/x where x is the volume of the smaller similar solid.

Directrix (directrix):

(2/1) ^ 3 = 240 /x Cross multiply (after you cube 2) and then solve for x. @samsteronbroadway

Directrix (directrix):